OpenStudy (anonymous):

will fan and medal! Can somebody help me solve this with an explanation please!! What is the APR of a payday loan for $1100 due in 5 days that charges a $165 fee?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

the borrower receives $1100 today, and must pay $1265 in 5 days

OpenStudy (perl):

how many 5 day periods are in a 365 day year?

OpenStudy (perl):

its a little more involved

OpenStudy (anonymous):

so 73?

OpenStudy (perl):

365/5 = 73 periods of 5 days in a year. 73 * $165 = $12,045 total interest in a year (if he kept that rate in 5 days)

OpenStudy (anonymous):

but then how would we find the percentage?

OpenStudy (perl):

12,045 / 1100 = 10.95 = 10.95 * 100 %

OpenStudy (perl):

1095% APR

OpenStudy (anonymous):

thank you so much i will refer to this every time i come across a question like this!!

OpenStudy (anonymous):

life saver like for reals!